OpenStudy (anonymous):

Peter is 4 years older than Dianne. In 6 years time, Peter will be 3 times as old as Dianne was 2 years ago. Find their present age.

OpenStudy (anonymous):

let x be Dianne's current age. y be Peter's current age. x+4=y. (Peter is 4 years older than Dianne) 3*(x-2) = y+6. (3 * Dianne's age 2 years ago, is equal to Peter's age in 6 years time) Now you have two equations and two unknowns. You can solve this and you should get that Dianne is 8 years old and Peter is 12 years old. so 2 years ago Dianne was 6. And in 6 years Peter will be 18. 6 * 3 = 18, so this is the right answer. It is hard figuring out how to write out the problem, but you just have to turn it into equations.